# Pushing Trade-Off Boundaries: Compact yet Effective Remote Sensing Change Detection

### v1

@@ 5 Conclusion
We propose a lightweight model, FlickCD, which maintains high performance while minimizing resource requirements. The model is primarily composed of convolutional operations and attention mechanisms. Specifically, we design the Enhanced Difference Module (EDM) to enhance bitemporal image difference information while reducing irrelevant noise. Additionally, we introduce two lightweight feature extraction modules, SWSA and EGSA, which effectively extract the local and global features while maintaining efficiency. To evaluate the effectiveness of FlickCD, we conducted extensive experiments on four datasets, analyzing its resource consumption and inference speed, along with ablation studies. The results consistently demonstrate our model’s low computational cost and high performance. This research provides new insights and methodologies for deploying change detection models on mobile devices in the future.

### v2

@@ 5 Conclusion
We propose a lightweight model, FlickCD, which maintains high performance while minimizing resource requirements. The model is primarily composed of convolution operations and attention mechanisms. Specifically, we design the Enhanced Difference Module (EDM) to enhance bi-temporal image difference information while reducing irrelevant noise. Additionally, we introduce two lightweight feature extraction modules, SWSA and EGSA, which effectively extract the local and global features while maintaining efficiency. To evaluate the effectiveness of FlickCD, we conducted extensive experiments on four datasets, analyzing its resource consumption and inference speed, along with ablation studies. The results consistently demonstrate our model’s low computational cost and high performance. This research provides new insights and methodologies for deploying change detection models on resource-constrained devices in the future.